OpenStudy (anonymous):

What is f(t) = y^2 + 9, what is f(t) when t = 4?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

I meant to say when

OpenStudy (amistre64):

4 = y^2 +9 -5 = y^2 y=+-isqrt(5)

OpenStudy (amistre64):

4 = y^2 +9 -5 = y^2 y=+-isqrt(5)

OpenStudy (amistre64):

nah, that aint it....

OpenStudy (amistre64):

can you restate the question to be clearer?